AUG011
Hb
Module HOB+1000, +5509 GB
- Good pointing fit
- Fringes found with Ke & Yg.
- Exp. started ok.
- 00:56 pcfshb ran out of disk space. Did not trigger log monitor alarm, and wasn't noticed until 02:00. No scans missed.
- 14:45 UT : disc pos ~170 GB behind pos, Rxmon restarted. (Ellen)
Ke
Module HOB+0096, +70 GB
- OK pointing fit
- Fringes found with Hb
- Exp. started ok
- 02:29 UT — 03:15 Ut. Sporadically high tsys in X band. Seems to have cleared. Possible RFI?
- I am witnessing very high X-band Tsys occasionally (Chris)
Yg
Module HOB+0114, +1 GB
- Good pointing fit
- Fringes found with Hb
- Exp. started ok.
~~DISCUSSION~~